# How Animals Dance (When You’re Not Looking)

### v1

@@
Abstract.
We present a keyframe-based framework for generating music-synchronized, choreography aware animal dance videos. Starting from a few keyframes representing distinct animal poses—generated via text-to-image prompting or GPT-4o— we formulate dance synthesis as a graph optimization problem: find the optimal keyframe structure that satisfies a specified choreography pattern of beats, which can be automatically estimated from a reference dance video. We also introduce an approach for mirrored pose image generation, essential for capturing symmetry in dance. In-between frames are synthesized using an video diffusion model. With as few as six input keyframes, our method can produce up to 30 second dance videos across a wide range of animals and music tracks. See our project page for more results: how-animals-dance.github.io.
dance video generation, diffusion models, vision for graphics
journal: TOGjournalyear: 2025ccs: Computing methodologies Computer vision
Figure 1. Starting from a small set of generated keyframes, e.g., a marmot in various poses, our method generates an animal dance video that follows a specified choreography pattern, extracted from a reference dance video.
1. Introduction
@@
Humans dance spontaneously to music—just picture a toddler cheerfully bouncing to the beats at a birthday party. Animals can dance too; Snowball the cockatoo (Fig. 2)—can perform up to $14$ distinct dance movements in response to different musical cues (Keehn et al., 2019). In fact, our animal friends are probably dancing all the time when we’re not looking. In this paper, we capture this hidden world of animal dance, and expose it for the first time to the human population (Fig. 1).
While we happen to be particularly obsessed with dancing animals, this paper introduces a new framework for generating music-synchronized, highly structured, 30 second+ long dance videos. Such capabilities are challenging for current state of the art generative models, most of which are limited to short clips of a few seconds, do not produced synchronized audio and video, and lack intuitive controls for long range motion. Beyond text prompting, most controls for video generation are fine-grained and operate on a single frame at a time, e.g., body pose, camera pose, motion brushes, etc. In contrast, we introduce choreography patterns as a new control for video generation. Specifically, we allow the user to specify a structured sequence of dance moves, or “beats”, e.g., A-B-A-B-C-D-A, where each letter corresponds to a particular move, and constrain the motion in the video to follow that choreography. Furthermore, we show how these choreography patterns can be automatically estimated from existing (human) dance videos.
A well-formed dance follows basic choreographic rules (Chen et al., 2021), which structure the movements to align with the rhythmic flow of the accompanying music, and often involve recurring patterns such as mirroring and repetition to help reinforce the musical structure (Kim et al., 2003, 2006). We leverage this inherent structure of dance to make the generation task more tractable. As input, we use a collection of initially generated keyframes, each representing a distinct pose. We then formulate the dance synthesis as a graph optimization problem, i.e., find the optimized walk path through the keyframes to satisfy a specified choreography pattern of beats. Each selected keyframe in the path is aligned to a musical beat. The final dance video is produced by synthesizing in-between frames between the keyframes using a generative video inbetweening model (Wang et al., 2024b, a) (Fig. 3).
Figure 2. Snowball the cockatoo: a real bird that can dance.
@@ 2. Prior Work
Music-driven generative dance synthesis. Earlier learning-based approaches developed neural networks that synthesize human dance motion directly from music input (Lee et al., 2018, 2019; Huang et al., 2020; Li et al., 2021; Zhang et al., 2022; Sun et al., 2020). Recent advances have shifted toward diffusion-based methods (Qi et al., 2023; Le et al., 2023; Tseng et al., 2023), which also focus primarily on generating skeletal motions from music. More recently, some works have begun exploring direct dance video generation using video diffusion models (Sun et al., 2020; Ruan et al., 2023). However, directly enforcing choreography structure within these frameworks remains challenging. In addition, these learning-based approaches typically require training data—an issue in our case, as dance videos featuring animals are extremely scarce.
Figure 3. System overview. Given a few initially generated keyframes as input, we generate mirrored counterparts, extract choreography pattern from a dance video, and optimize the keyframe structure accordingly. The final dance is synthesized by generating in-between frames with a video diffusion model and warped to the musical beats. Our method is highlighted in gray.
@@
Graph-based human motion synthesis. In contrast to learning-based approaches, graph-based frameworks (Kim et al., 2003, 2006; Ofli et al., 2008; Manfrè et al., 2016; Chen et al., 2021) synthesize new motions from an existing dance motion segments database, and cast the dance synthesis as a graph optimization problem: finding an optimal path in the constructed motion graph that aligns with the input music. For example, Kim et al. (Kim et al., 2003) introduced rhythmic and beat-based constraints to guide the path search, while more recent work ChoreoMaster (Chen et al., 2021) incorporated richer choreography rules, requiring not only structural alignment with music but also stylistic compatibility.
@@
Our approach follows this paradigm but differs in key ways. First, instead of relying on a motion capture database, we take a small set of keyframes of an animal or subject as input. We augment this set by generating mirrored pose images, creating a complete keyframe set for dance synthesis. The graph is constructed over these keyframes, and a video diffusion model is applied to generate realistic in-between frames along the optimized walk path, producing the final dance video. Second, while basic choreography rules can be inferred from the musical structure as done in  (Chen et al., 2021), different performers may interpret the same piece differently. As such, we propose a way to extract choreography patterns directly from a reference dance video and use it as the control.
@@
Anthropomorphic character animation. Given a reference image, character image animation, generates videos following a target human skeletal pose sequence. While existing methods (Hu, 2024; Hu et al., 2025) are primarily designed for human figures, recent work (Tan et al., 2024) extends to anthropomorphic characters by learning generalized motion representation. However, it remains favor human-like anatomy, generating animals with features like elongated limbs. They also struggle to generate high-fidelity videos from a single image when handling long and diverse sequences, such as a 30-second dance. In contrast, our method does not rely on per-frame human skeleton pose as guidance. Instead, we use a small set of keyframes generated from text prompt as a pose alternative, and synthesize intermediate frames using a generative inbetweening model. Since our approach is keyframe-based and choreographically structured, it is not constrained by video length and can scale to longer dance sequences.
3. Approach
@@
3.1. Choreography Pattern Labeling
@@
Problem definition. Given a 4/4 music track with a synchronized dance video, we begin by detecting the beat times =$\mathcal{B}=\{t_{0},t_{1},...,t_{N-1}\}$, assuming a total of $N$ beats, corresponding to $\frac{N}{4}$ bars. Based on the beat times, we construct a sequence of motion segments $\mathcal{S}=\{s_{0},s_{1},...,s_{\frac{N}{2}-1}\}$ where each segment $s_{i}$ spans from beat $t_{2i}$ to beat $t_{2i+1}$. Each bar thus yields two motion segments: one from the first to the second beat, and another from the third to the fourth beat, aligning with the 4/4 music structure where major movements typically begin on accented beats and end on weak ones, whereas transitions occur across weak-to-accented intervals. The “choreo pattern” labeling task outputs a sequence of labels =$\mathcal{L}=\{l_{0},l_{1},...,l_{\frac{N}{2}-1}\}$, e.g., A-A’-B-C-D-D, where each $l_{i}$ corresponds to motion segment $s_{i}$. Distinct motions receive different labels, identical motions share the same one, and mirrored motions are indicated by prime-labeled counterparts (e.g., A and A’).
Motion segments quantization. We formulate motion segment sequence labeling as a quantization problem: clustering similar motion segments and assigning each a cluster ID as its label. Each segment $s_{i}$ of length $T_{i}$ is represented by the SMPL-X (Pavlakos et al., 2019) pose sequence recovered from the video: $s_{i}=\{(\theta_{t_{i}}\in\mathbb{R}^{3\times(J+1)},\tau_{t_{i}}\in\mathbb{R}^{3})\}_{t_{i}=0}^{T_{i}}$, where $\theta_{t_{i}}$ contains per-joint axis-angle rotation for $J=21$ body joints in addition to a joint for global rotation (the 00-th joint), and $\tau_{t_{i}}$ denotes the global translation in 3D space.
@@
(1) $d_{\theta}(\theta_{1},\theta_{2})=\frac{1}{J}\sum_{j=0}^{J}||\log(R(\theta_{1}^{j})^{T}R(\theta_{2}^{j}))||_{F}$
@@
(2) $\mathcal{C}(\mathcal{S};\text{DTW}_{d_{\theta}},\epsilon_{\theta})\rightarrow\{\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{C}\}$
@@
(3) $\theta^{\prime j}=\mathcal{F}(\theta^{\pi(j)})\quad\text{for }j=0,\dots,J$
@@
Two clusters $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ are considered mirrored if there exists at least one pair of segments $s_{i}\in\mathcal{C}_{a},s_{j}\in\mathcal{C}_{b}$, such that the mirrored version of $s_{i}$, denoted by $s^{\prime}_{i}=\{\theta^{\prime}_{t_{i}}\}_{t_{i}=0}^{T_{i}}$, is similar to $s_{j}$ under dynamic time warping: $\text{DTW}_{d_{\theta}}(s^{\prime}_{i},s_{j})<\epsilon_{\theta^{\prime}}$. The resulting set of mirrored cluster pairs is: =$\mathcal{M}=\{(\mathcal{C}_{a},\mathcal{C}_{b})|\ \exists s_{i}\in\mathcal{C}_{a},s_{j}\in\mathcal{C}_{b},\text{DTW}_{d_{\theta}}(s^{\prime}_{i},s_{j})<\epsilon_{\theta^{\prime}}\}$.
Mirrored motion directions within a cluster. For clusters without a mirrored counterpart, we further check whether they can be internally partitioned into two directionally mirrored groups. We first extract the overall motion direction $\vec{d}_{s_{i}}$ of each motion segment $s_{i}$ using its global translation: $\vec{d}_{s_{i}}=(\tau_{t_{i}}^{T_{i}}-\tau_{t_{i}}^{0})/||\tau_{t_{i}}^{T_{i}}-\tau_{t_{i}}^{0}||$, where $\tau_{t_{i}}^{0}$and $\tau_{t_{i}}^{T_{i}}$ denote the segment’s start and end positions, respectively.
@@
3.2. Mirrored Pose Image Generation
To augment the keyframe set with mirrored counterparts, we generate visually consistent keyframe pairs for each input pose. This process (Fig. 4) involves fine-tuning a text-to-image model with ControlNet, generating mirrored edge maps, and re-generating the original keyframes for consistency. In the end, we get a complete set of consistent keyframes =$\mathcal{I}=\{I_{0},\dots,I_{K-1},I^{\prime}_{0},\dots,I^{\prime}_{K-1}\}$, where $I^{\prime}_{k}$ is the mirrored version of $I_{k}$.
Fine-tuning. We fine-tune a pretrained text-to-image model on the input keyframes set, overfitting it to capture the appearance of the specific input subject instance and the background. To provide structural guidance, we incorporate ControlNet (Zhang et al., 2023) using the canny edge maps extracted from the input images as a conditional input. We use the prompt format: “A photo of [V] [subject class name] dancing [background description].”, where [V] is a unique token for identifying the subject instance rather than class. The placeholders [subject class name] and [background description] are replaced with the actual class name of the subject and background description. For example, “A photo of [V] marmot dancing with alpine landscape as background.”
Mirrored edge generation. To generate mirrored pose images, We first extract the subject mask using SAM (Kirillov et al., 2023). We then construct a unified background canny edge map by inpainting and stitching the background edges from all input keyframes. For each keyframe, we extract the subject’s edge map and horizontally flip it to create a mirrored subject edge map. This flipped edge map is then composited with the shared background edge map to generate a full mirrored edge map. which is used as input to the fine-tuned model to generate the corresponding mirrored image.
Keyframes re-generation. The original keyframes may contain slight inconsistencies in the background due to generation instability. Additionally, the fine-tuned model may introduce subtle color shifts during inference. To ensure visual consistency among the augmented keyframes set, we regenerate the original keyframes using the same model and shared background edge map (see Fig. 5 for an example).
Figure 4. Mirrored pose generation. We fine-tune a text-to-image model with ControlNet using the canny edges extracted from each keyframe as conditioning. During inference, mirrored pose images are generated by flipping only the subject edges and using an inpainted background edge composed from the keyframe set.
@@ 3.2. Mirrored Pose Image Generation
(b) re-generated keyframes
Figure 5. Improving visual consistency by re-generating keyframes with shared background edges.
3.3. Choreography Pattern Driven Dance Synthesis
Given the augmented keyframe set =$\mathcal{I}=\{I_{0},\dots,I_{K-1},I^{\prime}_{0},\dots,I^{\prime}_{K-1}\}$, where $I_{k}^{\prime}$ denotes the mirrored counterpart of keyframe $I_{k}$, and the choreography pattern label sequence =$\mathcal{L}=\{l_{0},l_{1},...,l_{\frac{N}{2}-1}\}$, the goal is to find an optimal walk path $\mathcal{P}=\{I_{p_{0}},I_{p_{1}},\dots,I_{p_{N-1}}\}$ through the keyframe set, and the $i$-th keyframe $I_{p_{i}}$ in the path corresponds to the $i$-th beat. We then apply video diffusion model to generate in-between frames, finally producing the final dance video.

### v2

@@
Abstract
We present a framework for generating music-synchronized, choreography aware animal dance videos. Our framework introduces choreography patterns—structured sequences of motion beats that define the long-range structure of a dance—as a novel high-level control signal for dance video generation. These patterns can be automatically estimated from human dance videos. Starting from a few keyframes representing distinct animal poses, generated via text-to-image prompting or GPT-4o, we formulate dance synthesis as a graph optimization problem that seeks the optimal keyframe structure to satisfy a specified choreography pattern of beats. We also introduce an approach for mirrored pose image generation, essential for capturing symmetry in dance. In-between frames are synthesized using an video diffusion model. With as few as six input keyframes, our method can produce up to 30 seconds dance videos across a wide range of animals and music tracks.
1 Introduction
@@
Humans dance spontaneously to music—just picture a toddler cheerfully bouncing to the beats at a birthday party. Animals can dance too; Snowball the cockatoo can perform up to $14$ distinct dance movements in response to different musical cues [keehn2019spontaneity]. In fact, our animal friends are probably dancing all the time when we’re not looking. In this paper, we capture this hidden world of animal dance, and expose it for the first time to the human eyes.
While we happen to be particularly obsessed with dancing animals, this paper introduces a new framework for generating music-synchronized, highly structured, up to 30 seconds long dance videos. Such capabilities are challenging for current state of the art generative models [blattmann2023stable, bar2024lumiere, zeng2024make, yang2024cogvideox], most of which are limited to short clips of a few seconds, do not produced synchronized audio and video, and lack intuitive controls for long range motion. Beyond text prompting, most controls for video generation are fine-grained and operate on a single frame at a time [wang2024motionctrl, geng2024motionprompting], e.g., body pose, camera pose, motion brushes, etc. In contrast, we introduce choreography patterns as a new control for video generation. Specifically, we allow the user to specify a structured sequence of dance moves, or “beats”, e.g., A-B-A-B-C-D-A, where each letter corresponds to a particular move, and constrain the motion in the video to follow that choreography. Furthermore, we show how these choreography patterns can be automatically estimated from existing (human) dance videos.
Our use of choreography patterns as a control is inspired by how real dances are organized. A well-formed dance follows basic choreographic rules [choreomaster2021], which structure the movements to align with the rhythmic flow of the accompanying music, and often involve recurring patterns such as mirroring and repetition to help reinforce the musical structure [kim2003rhythmic, kim2006making]. We leverage this inherent structure of dance to make the generation task more tractable. As input, we use a collection of initially generated keyframes, each representing a distinct pose. We then formulate the dance synthesis as a graph optimization problem, i.e., find the optimal walk path through the keyframes where the underlying motion satisfy a specified choreography pattern of beats. Each selected keyframe in the path is aligned to a musical beat. The final dance video is produced by synthesizing in-between frames between the keyframes using a generative video inbetweening model [wang2024generative, wang2024framer] (Fig. 1).
Figure 1: System overview. Given a few initially generated keyframes as input, we generate mirrored counterparts, extract choreography pattern from a dance video, and optimize the keyframe structure accordingly. The final dance is synthesized by generating in-between frames with a video diffusion model and warped to the musical beats. Our method is highlighted in gray.
@@ 2 Prior Work
Music-driven generative dance synthesis. Earlier learning-based approaches developed neural networks that synthesize human dance motion directly from music input [lee2018listen, lee2019dancing, huang2020dance, li2021ai, zhang2022music, sun2020deepdance]. Recent advances have shifted toward diffusion-based methods [qi2023diffdance, le2023controllable, tseng2023edge], which also focus primarily on generating skeletal motions from music. More recently, some works have begun exploring direct dance video generation using video diffusion models [sun2020deepdance, ruan2023mm, hong2025musicinfuser]. However, directly enforcing choreography structure within these frameworks remains challenging. In addition, these learning-based approaches typically require training data—an issue in our case, as dance videos featuring animals are extremely scarce.
@@
Graph-based human dance motion synthesis. In contrast to learning-based approaches, graph-based frameworks [kim2003rhythmic, kim2006making, ofli2008audio, manfre2016automatic, choreomaster2021] synthesize new motions from an existing dance motion segments database, and cast the dance synthesis as a graph optimization problem: finding an optimal path in the constructed motion graph that aligns with the input music. For example, Kim et al. [kim2003rhythmic] introduced rhythmic and beat-based constraints to guide the path search, while more recent work ChoreoMaster [choreomaster2021] incorporated richer choreography rules, requiring not only structural alignment with music but also stylistic compatibility.
@@
Our approach follows this paradigm but differs in key ways. First, instead of relying on a motion capture database, we take a small set of keyframes of an animal or subject as input. We augment this set by generating mirrored pose images, creating a complete keyframe set for dance synthesis. The graph is constructed over these keyframes, and a video diffusion model is applied to generate realistic in-between frames along the optimized walk path, producing the final dance video. Second, while basic choreography rules can be inferred from the musical structure as done in  [choreomaster2021], different performers may interpret the same piece differently. As such, we propose a way to extract choreography patterns directly from a reference dance video and use it as the control.
@@
Anthropomorphic character animation. Given a reference image, character image animation, generates videos following a per-frame target human skeletal pose sequence. While existing methods [hu2024animate, hu2025animate] are primarily designed for human figures, recent work [tan2024animate] extends to anthropomorphic characters by learning generalized motion representation. However, it still favor human-like anatomy, often producing animals with features like elongated limbs and human-style body proportions. They also struggle to generate high-fidelity videos from a single image when handling long and diverse sequences, such as a 30-second dance. In contrast, our method does not rely on per-frame human skeleton pose as guidance and uses choreography pattern as higher-level control, letting the video diffusion model generate in-between motions so that the final dance follows the choreographic structure, not human motion itself.
3 Approach
@@
3.1 Choreography Pattern Labeling
@@
Problem definition. Given a 4/4 music track with a synchronized dance video, we begin by detecting the beat times $\mathcal{B}=\{t_{0},t_{1},...,t_{N-1}\}$, assuming a total of $N$ beats, corresponding to $\frac{N}{4}$ bars. Based on the beat times, we construct a sequence of motion segments $\mathcal{S}=\{s_{0},s_{1},...,s_{\frac{N}{2}-1}\}$ where each segment $s_{i}$ spans from beat $t_{2i}$ to beat $t_{2i+1}$. Each bar thus yields two motion segments: one from the first to the second beat, and another from the third to the fourth beat, aligning with the 4/4 music structure where major movements typically begin on accented beats and end on weak ones, whereas transitions occur across weak-to-accented intervals. The “choreo pattern” labeling task outputs a sequence of labels $\mathcal{L}=\{l_{0},l_{1},...,l_{\frac{N}{2}-1}\}$, e.g., A-A’-B-C-D-D, where each $l_{i}$ corresponds to motion segment $s_{i}$. Distinct motions receive different labels, identical motions share the same one, and mirrored motions are indicated by prime-labeled counterparts (e.g., A and A’).
Motion segments quantization. We formulate motion segment sequence labeling as a quantization problem: clustering similar motion segments and assigning each a cluster ID as its label. Each segment $s_{i}$ of length $T_{i}$ is represented by the SMPL-X [SMPL-X:2019] pose sequence recovered from the video: $s_{i}=\{(\theta_{t_{i}}\in\mathbb{R}^{3\times(J+1)},\tau_{t_{i}}\in\mathbb{R}^{3})\}_{t_{i}=0}^{T_{i}}$, where $\theta_{t_{i}}$ contains per-joint axis-angle rotation for $J=21$ body joints in addition to a joint for global rotation (the 0-th joint), and $\tau_{t_{i}}$ denotes the global translation in 3D space.
@@
[…]
@@
$\mathcal{C}(\mathcal{S};\text{DTW}_{d_{\theta}},\epsilon_{\theta})\rightarrow\{\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{C}\}$ (2)
@@
$\theta^{\prime j}=\mathcal{F}(\theta^{\pi(j)})\quad\text{for }j=0,\dots,J$ (3)
@@
Two clusters $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ are considered mirrored if there exists at least one pair of segments $s_{i}\in\mathcal{C}_{a},s_{j}\in\mathcal{C}_{b}$, such that the mirrored version of $s_{i}$, denoted by $s^{\prime}_{i}=\{\theta^{\prime}_{t_{i}}\}_{t_{i}=0}^{T_{i}}$, is similar to $s_{j}$ under dynamic time warping: $\text{DTW}_{d_{\theta}}(s^{\prime}_{i},s_{j})<\epsilon_{\theta^{\prime}}$. The resulting set of mirrored cluster pairs is: $\mathcal{M}=\{(\mathcal{C}_{a},\mathcal{C}_{b})|\ \exists s_{i}\in\mathcal{C}_{a},s_{j}\in\mathcal{C}_{b},\text{DTW}_{d_{\theta}}(s^{\prime}_{i},s_{j})<\epsilon_{\theta^{\prime}}\}$.
[…]
@@
3.2 Mirrored Pose Image Generation
Figure 2: Mirrored pose generation. We fine-tune a text-to-image model with ControlNet using the canny edges extracted from each keyframe as conditioning. During inference, mirrored pose images are generated by flipping only the subject edges and using an inpainted background edge composed from the keyframe set.
To augment the keyframe set with mirrored counterparts, we generate visually consistent keyframe pairs for each input pose. This process (Fig. 2) involves fine-tuning a text-to-image model with ControlNet, generating mirrored edge maps, and re-generating the original keyframes for consistency. In the end, we get a complete set of consistent keyframes $\mathcal{I}=\{I_{0},\dots,I_{K-1},I^{\prime}_{0},\dots,I^{\prime}_{K-1}\}$, where $I^{\prime}_{k}$ is the mirrored version of $I_{k}$.
Fine-tuning. We fine-tune a pretrained text-to-image model on the input keyframes set, overfitting it to capture the appearance of the specific input subject instance and the background. To provide structural guidance, we incorporate ControlNet [zhang2023adding] using the canny edge maps extracted from the input images as a conditional input. We use the prompt format: “A photo of [V] [subject class name] dancing [background description].”, where [V] is a unique token for identifying the subject instance rather than class. The placeholders [subject class name] and [background description] are replaced with the actual class name of the subject and background description. For example, “A photo of [V] marmot dancing with alpine landscape as background.”
Mirrored edge generation. To generate mirrored pose images, We first extract the subject mask using SAM [kirillov2023segany]. We then construct a unified background canny edge map by inpainting and stitching the background edges from all input keyframes. For each keyframe, we extract the subject’s edge map and horizontally flip it to create a mirrored subject edge map. This flipped edge map is then composited with the shared background edge map to generate a full mirrored edge map. which is used as input to the fine-tuned model to generate the corresponding mirrored image.
Keyframes re-generation.
@@ 3.2 Mirrored Pose Image Generation
(b) refined keyframes
Figure 3: Improving visual consistency by re-generating keyframes with shared background edges.
The original keyframes may contain slight inconsistencies in the background due to generation instability. Additionally, the fine-tuned model may introduce subtle color shifts during inference. To ensure visual consistency among the augmented keyframes set, we regenerate the original keyframes using the same model and shared background edge map (see Fig. 3 for an example).
3.3 Choreography Pattern Driven Dance Synthesis
Given the augmented keyframe set $\mathcal{I}=\{I_{0},\dots,I_{K-1},I^{\prime}_{0},\dots,I^{\prime}_{K-1}\}$, where $I_{k}^{\prime}$ denotes the mirrored counterpart of keyframe $I_{k}$, and the choreography pattern label sequence $\mathcal{L}=\{l_{0},l_{1},...,l_{\frac{N}{2}-1}\}$, the goal is to find an optimal walk path $\mathcal{P}=\{I_{p_{0}},I_{p_{1}},\dots,I_{p_{N-1}}\}$ through the keyframe set, and the $i$-th keyframe $I_{p_{i}}$ in the path corresponds to the $i$-th beat. We then apply video diffusion model to generate in-between frames, finally producing the final dance video.